OpenStudy (anonymous):

Find the value of the polynomial (2y - 3)3 when y=9

OpenStudy (campbell_st):

(2 x 9 -3)^3 = (15)^3 you can finish it

OpenStudy (anonymous):

value = [2(9) - 3]^3 = [18 - 3]^3 = 15^3 = 3375